4. Consider the equilibrium reaction:
2 SO2 (g) + O2(g) ⇌ 2SO3(g)
In a particular experiment a student added 0.0870 moles of SO2 and 0.0780 moles of O2 to a 0.750-L
container. At equilibrium, the equilibrium concentration of O2 was 0.065 molar. What is the value of
the equilibrium constant K?

5. Consider the reaction
2 A(aq) ⇌ B(aq) + C(aq)
At 77oC, the value of K is 264. A 0.750 mole sample of A is added to a 1.5 L container. What are the
concentrations of all species when equilibrium is established?

6. The reaction:
CO(g) + Cl2 (g) ⇌ COCl2(g)
has an equilibrium constant of 1.75 at 300oC. A 1.0 L flask was filled with 1.16 mol each of carbon
monoxide and chlorine. What are the concentrations of all substances present when the system
reaches equilibrium?
1
Expert's answer
2020-10-16T08:57:15-0400

Solution.

1.

"C(SO2) = \frac{0.0870}{0.75} = 0.116 \ M"

"C(O2) = \frac{0.0780}{0.75} = 0.104 \ M"

"C(new \ O2) = \frac{0.065}{0.75} = 0.087 M"

"C(SO3) = 2*(0.104-0.087) = 0.034"

"C(new SO2) = 0.116-(2\times0.017) = 0.082 M"

"K = \frac{C(SO3)^2}{C(NewSO2)^2 \times C(newO2)} = 1.98"

2.

Solution.

"264 = \frac{x^2}{0.5-2\times x}"

x = 0.24

C(A) = 0.02 M

C(B) = 0.24

C(C) = 0.24

3.

Solution.

"1.75 = \frac{x^2}{(1.16-x)^2}"

x = 0.66

C(COCl2) = 0.66 M

C(CO) = 0.5 M

C(Cl2) = 0.5 M

Answer:

1.

K = 1.98

2.

C(A) = 0.02 M

C(B) = 0.24

C(C) = 0.24

3.

C(COCl2) = 0.66 M

C(CO) = 0.5 M

C(Cl2) = 0.5 M